OpenStudy (anonymous):

prove the A=(6,2),B=(3,3) and c=(0,4) are on a straight line

OpenStudy (kira_yamato):

BA vector = <6-3 , 2-3> = <3 , -1> CA vector = <6-0 , 2 - 4> = <6 , -2> = 2*<3 , -1> = 2(BA vector) Therefore, A, B, C are collinear (QED)

OpenStudy (anonymous):

thank you